OpenStudy (anonymous):

you bought a new fish aquarium with the dimensions of 55cm x 100cm x 80. what volume of water should you put in? show work

OpenStudy (anonymous):

V=l*w*h 55*100*80= 440000cm^3 * 1m^3/1000000cm^3 *1000L/1m^3 =440L Of course you would never fill an aquarium to the brim, so 420L should be enough to blaze it.

OpenStudy (anonymous):

please explain why you would do 1m^3/1000000cm^3*1000L/1m^3

OpenStudy (anonymous):

Not neccessary but there is 1*10^6 cm3 in a m3, and 1000Liters in a m3. 440000cm^3 is fine in terms of volume, but who fills an aquarium measuring in cm^3?

OpenStudy (anonymous):

so you convert the unit and then you what??

OpenStudy (anonymous):

Im sorry this is really confusing for me

OpenStudy (anonymous):

Its just unit conversion from cm^3 to liters. 440L=440000cm^3 of water. If the question is asking how much you should fill the aquarium, you'd typically not fill it to the top but knowing how school is they probably just want the volume.

OpenStudy (anonymous):

ok so 440L would be the volume of how much water can be filled into the tank from all the way to the top

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok thank you nukka